OpenStudy (jkbo):

Please help The slope of x = 1 is undefined. A.True B.False

jimthompson5910 (jim_thompson5910):

hint: x = 1 is a vertical line through x = 1 on the x axis

OpenStudy (jkbo):

so its false @jim_thompson5910

jimthompson5910 (jim_thompson5910):

are you sure?

OpenStudy (jkbo):

yes @jim_thompson5910

jimthompson5910 (jim_thompson5910):

well horizontal lines have a slope of 0 vertical lines have an undefined slope

OpenStudy (jkbo):

o ok so it false

OpenStudy (jkbo):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

which one is it? true or false?

OpenStudy (jkbo):

its false @jim_thompson5910

jimthompson5910 (jim_thompson5910):

horizontal lines have a slope of 0 vertical lines have an undefined slope what type of line is x = 1 again? vertical or horizontal?

OpenStudy (jkbo):

horizontal @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no that would be things like y = 1 or y = 10, etc

OpenStudy (jkbo):

ok then its verical

jimthompson5910 (jim_thompson5910):

yes so because it's vertical, it has an undefined slope

OpenStudy (jkbo):

that mean it true? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes the answer is true

OpenStudy (jkbo):

thank you can u help me with one more @jim_thompson5910 Write the slope-intercept form of the equation for the line A.y = 3x - 1 B.y = 1/3 x + 1 C.y = -3x - 1 D.y = 1/3 x - 1

jimthompson5910 (jim_thompson5910):

look at the graph what is the slope of this line?

OpenStudy (jkbo):

1/3 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

how are you getting 1/3?

OpenStudy (jkbo):

sorry 9/2 rise over run @jim_thompson5910

jimthompson5910 (jim_thompson5910):

getting closer, but no

OpenStudy (jkbo):

Im not sure then @jim_thompson5910

jimthompson5910 (jim_thompson5910):

start at the first point, then what is the rise? what is the run? to get to the next point

OpenStudy (jkbo):

Rise ;4 run 2 thats what I got @jim_thompson5910

jimthompson5910 (jim_thompson5910):

it should be rise = 6 run = 2 so the slope is what?

OpenStudy (jkbo):

3/1 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

or just 3

jimthompson5910 (jim_thompson5910):

now what is the y-intercept

OpenStudy (jkbo):

1 right? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

where does it cross the y axis?

OpenStudy (jkbo):

i'm not sure @jim_thompson5910

jimthompson5910 (jim_thompson5910):

look at the graph

jimthompson5910 (jim_thompson5910):

where does the diagonal line cross the y axis?

OpenStudy (jkbo):

it crosses at 1 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

look again

jimthompson5910 (jim_thompson5910):

the y axis is the vertical axis

OpenStudy (jkbo):

I cant find it @jim_thompson5910

jimthompson5910 (jim_thompson5910):

see attached (look at the red dot)

OpenStudy (jkbo):

o ok -1 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

so that's the y-intercept

OpenStudy (jkbo):

o so its y=3x-1 A.

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (jkbo):

thanks

jimthompson5910 (jim_thompson5910):

np